# arXiv:2603.00989v2[cs.SE] 03 Mar 2026

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used in software engineering to help developers generate, complete, translate, and fix code. These tools improve productivity and speed up the development process. However, their growing use has raised concerns about environmental sustainability. Most existing studies focus on the high energy consumption and carbon emissions of model training and inference. In contrast, much less attention has been given to the sustainability of the code generated by these models. The efficiency of generated code affects the long-term environmental impact of software systems. Inefficient code can increase CPU usage, memory consumption, execution time, and overall energy use during deployment and operation. As LLM-generated code becomes more common in real-world projects, even small inefficiencies can lead to high environmental costs over time. This paper examines existing research on the sustainability of code generated by LLMs. We conduct a systematic literature review to analyze selected primary studies and investigate the extent to which LLMs are capable of producing sustainable code. In addition, we examine how sustainability is defined and measured in this context, including the metrics and evaluation strategies used to assess energy efficiency and resource usage. We also explore whether techniques such as fine-tuning and prompt engineering influence the sustainability of generated code. Through a structured analysis of the selected studies, we categorize research efforts based on their methodological approaches, evaluation practices, and experimental settings. The findings indicate that research in this area remains relatively limited and fragmented, with no widely accepted framework for measuring or benchmarking the sustainability of LLM-generated code. These observations highlight the need for clearer definitions, standardized evaluation methods, and systematic research to support environmentally friendly AI-assisted software engineering.
@@ I Introduction
Large Language Models (LLMs) are advanced AI systems trained on large corpora of data using transformer-based architectures [131]. These models are capable of understanding, reasoning over, and generating human-like language, enabling their use across a wide range of applications[150]. In recent years, LLMs have gained significant attention in the software engineering domain, where they support and automate various programming tasks. One of the most prominent applications is code generation, in which an LLM produces executable source code from natural language instructions. Closely related tasks include code completion (predicting missing or subsequent lines of code), code translation (converting source code between programming languages), bug fixing (detecting and correcting programming errors), and documentation generation (producing natural language explanations of code functionality). These capabilities are now embedded in widely used AI-assisted development tools, such as GitHub Copilot [43], CodeGeeX [144], and Amazon CodeWhisperer [8], which leverage state-of-the-art code LLMs to improve developer productivity and streamline the software development process.
LLMs have achieved tremendous success in Code generation [52, 104]; however, there is a growing concern about the impact of software development on the environment. Training and deploying LLMs incurs significant environmental costs, including substantial CO2 emissions and water usage [34]. According to [86], training LLaMA 3.1 with 8 billion parameters generated approximately 420 tCO2e, which is equivalent to the emissions from 83 years of electricity usage by a single U.S. household. The process also consumed 2,769 kiloliters of water, roughly equal to 24.5 years of water usage by an average American, and this is only for the training phase. Once deployed, these models continue to consume energy as users interact with them. Energy usage during inference has increased rapidly [34], and total emissions depend on how frequently the model is used. For example, if ChatGPT receives 100 million queries per day [34], and each query consumes about 0.002 kWh of energy, the total daily energy consumption would be approximately 0.2 gigawatt-hours (GWh) [129].
@@
The remainder of this paper is organized as follows. Section II provides background information on LLMs and code generation. Section III describes the research methodology. Section IV provides insights into the extracted data, while Section V presents the analysis and results from the extracted data. Section VI presents the future research in the area. Section VII outlines the validity threats and limitations. Finally, Section VIII presents concluding remarks.
@@ II-A Large Language Models
Architecturally, an LLM is a stack (concatenation) of Transformer blocks [131]. To pass the text data in the transformer, each input token (word) $x_{t}$ is first mapped to a continuous embedding [7] $e_{t}\;=\;E(x_{t})\in\mathbb{R}^{d}$ where $E$ is the embedding matrix and $d$ is the hidden size. A Transformer block takes the current hidden states and applies self-attention followed by a position-wise feed-forward network (with residual connections and layer normalization omitted here for brevity). This combination of self-attention and feed-forward layers, together with residual connections and normalization, forms the core building block of the Transformer architecture used in LLMs. For more information on transformer architectures and their underlying mechanisms, readers are referred to [131]
@@ II-A5 Fine-tuning and Prompt Engineering
TABLE I: Research Questions
RQ Description
RQ2 What metrics or parameters are used to evaluate sustainability?
@@ III-B Research Questions
RQ1 is motivated by the need to evaluate how well LLMs can generate code that meets sustainability goals, such as reducing energy consumption, lowering computational overhead, and minimizing resource usage. It examines whether sustainability is an inherent property of LLM-generated code or the result of deliberate design choices. The scope includes code generated across different application domains and programming languages. Addressing RQ1 will help researchers identify which LLMs produce more sustainable code and understand the factors that contribute to this outcome.
RQ2 aims to identify and categorize the measures used to evaluate the sustainability of LLM-generated code. These may include metrics such as energy consumption, carbon footprint, memory usage, execution time, and other efficiency indicators. Examining these metrics will show how sustainability is defined and assessed in existing studies. The findings will also support the development of standardized evaluation methods, enabling more consistent comparisons across future research.
RQ3 focuses on identifying the resources used to assess the sustainability of LLM-generated code, including benchmarks, datasets, simulation frameworks, and performance analysis tools. It examines whether these resources are specifically designed for sustainability evaluation or adapted from general-purpose code generation benchmarks. The findings will help researchers choose appropriate resources for experimentation and highlight gaps where new sustainability-oriented datasets or tools are needed.
RQ4 examines how model adaptation techniques affect the sustainability of generated code, including prompt engineering, instruction tuning, and domain-specific fine-tuning, etc. It focuses on the direct impacts of these techniques on LLM-generated code, such as reduced runtime and lower energy consumption. The findings will help establish best practices for designing prompts and fine-tuning methods that enhance sustainability without reducing functional correctness.
@@ III-C Search Strategy
Guided by the RQs, we adopt an iterative search strategy. We repeatedly query digital libraries using Boolean combinations of relevant keywords, and after each iteration, we refine the keywords to improve coverage. This process continues until the final search results include all publications from a manually identified benchmark set.
@@ III-C1 Search String
Our search term was structured around two main aspects of the topic. The first aspect relates to LLMs, code generation, and associated techniques such as fine-tuning. The second aspect focuses on sustainability and energy efficiency. Combining these aspects allowed us to capture the broadest possible set of studies relevant to sustainable code generation using LLMs. The final search term was formulated by linking keywords from both aspects to ensure that retrieved publications addressed concepts from each domain.
We refined the search string iteratively by evaluating the results it produced. To guide this process, we manually compiled a benchmark set of thirteen (13) relevant research publications from multiple sources, including Google Search, Google Scholar, and arXiv. In the first iteration, a comparison between the benchmark set and the search results showed that only 10 out of the 13 publications were retrieved. To capture the three missing publications, we added additional terms such as “LLM-generated code,” “energy efficiency,” and “efficient code” to the search string. After these additions, the revised search string successfully returned a result set containing the entire benchmark set. This final search string was adopted for the remainder of the study and is presented in Table III.
III-C2 Search Sources
We searched for relevant publications in five databases: IEEE Xplore, Compendex, Inspec, Scopus, and the ACM Digital Library. The retrieved publications were managed using an Excel spreadsheet. A detailed breakdown of the search process for each database, along with the corresponding number of publications, is presented in Fig. 1.
@@ III-C2 Search Sources
TABLE IV: Data Extraction Form
Category Data Items Description
Code-Related Task The primary coding task addressed (e.g., code generation, completion, etc.
Model Adaptation Fine-Tuning Details If the study is performing sustainability-aware fine-tuning.
III-D Inclusion and Exclusion Criteria
To ensure that only the most relevant and high-quality studies were considered in this systematic literature review, we applied a predefined set of inclusion and exclusion criteria, as summarized in Table II. In order to qualify as a primary study, the research work must meet all inclusion criteria and must not meet any exclusion criteria.
We included studies that used or evaluated LLMs for code generation, addressed sustainability aspects such as energy efficiency or resource usage, and provided empirical evaluation results. Studies were excluded if they did not involve an LLM, lacked sustainability analysis, focused on unrelated tasks, or had no accessible full text in English. The inclusion and exclusion criteria were applied at multiple stages of the literature review process, including the initial filtering of primary studies and the full-text screening phase. They were also employed during the snowballing process, which resulted in the addition of 4 more primary studies.
Although several studies claim efficiency improvements, many of them do not provide empirical evaluation of energy consumption or energy efficiency. Since our objective is to examine the sustainability of LLM-generated code based on measurable evidence, we refined our exclusion criteria to remove studies that do not include any form of energy consumption analysis. As a result, only studies that provide explicit energy-related evaluation were considered in the final selection. This refinement reduced the number of eligible studies to 19. Moreover, to ensure comprehensive coverage of this emerging research area, we also included relevant pre-prints from arXiv, as sustainability-focused research on LLM-generated code is still developing and may not yet be fully represented in peer-reviewed venues.
@@ III-D Inclusion and Exclusion Criteria
TABLE V: Overview of the evaluated models and their parameter-scale categories.
@@ III-E Study Selection Procedure
We then applied the refined inclusion criteria independently to the entire set of retrieved publications, identifying 170 studies for full-text review. To maintain consistency, we used the same pilot process with Fleiss’ Kappa measurement during this stage. After carefully reading the full texts, we found that only 15 studies fully met the refined criteria and were included as primary studies. We chose not to apply additional quality assessment filters, as the final pool of studies was already small and the research area is still emerging. This approach allowed us to remain as inclusive as possible while ensuring relevance to our research focus.
We conducted a snowballing [66] search to identify additional relevant primary studies. Snowballing involves reviewing the reference lists of papers (backward snowballing) and the papers that cite them (forward snowballing) to find further studies. In this study, both forward and backward snowballing were applied to the 15 retrieved primary studies. Google Scholar was used to obtain the citations for each paper, and the same selection criteria were applied to both their references and citations. This process yielded 4 additional relevant papers, bringing the final pool of primary studies to 19. Figure 1 shows how the initial 2,015 publications retrieved from the databases were reduced to these 19 primary studies, which were then used for data extraction.
III-F Data Extraction Form
We created a structured data extraction form to collect information from each primary study consistently and accurately, focusing on our research questions (RQ1–RQ4). First, three researchers prepared a draft extraction form based on a small sample of studies. We then met to combine these drafts into the final extraction form shown in Table IV. The categories in the form were designed to capture the main differences found in research on sustainable code generation using LLMs.
@@ III-F Data Extraction Form
IV Data Extraction
We present the analysis of data collected from the final pool of 19 primary studies. This section outlines the extracted data, as summarized in Table IV, providing a structured basis for analysis and enabling comparisons across all studies. The information is then used in Section LABEL:subsec:results_rqs to synthesize findings and address RQ1–RQ4.
IV-A General
IV-A1 Context of Study
We first categorized the selected studies based on the environment in which the research was undertaken, distinguishing between academic and industrial settings. A study was labelled as academic if all contributing authors were affiliated with universities or research institutions. A study was labelled as industrial if the research involved participation from industry, identified by the presence of at least one author with an industrial affiliation. The distribution of studies across these contexts is summarized in Table VI. Out of the 19 studies, 14 were conducted in an academic setting, while only 5 were performed in an industrial setting.
@@ IV-A1 Context of Study
IV-A2 Code-Related Task
We categorized the selected studies based on the code-related tasks they addressed. Code-related tasks refer to code generation, code completion, code optimization, and code refactoring. As several studies addressed more than one task (e.g., combining code generation with optimization or refactoring), a multi-label classification approach was adopted; therefore, a single study may appear in multiple task categories. This approach enables a more accurate representation of the scope of each study and avoids oversimplifying contributions that span multiple tasks.
@@ IV-A2 Code-Related Task
IV-A3 Code Application Domain
@@
General-Purpose Software Development [96], [110], [127], [22], [27], [60], [25], [130], [129], [64], [101], [11], [97], [116], [29], [12], [10] 17
@@
IV-B Model
IV-B1 Evaluated Models and Model Size Categories
We first present the list of models considered in each study in Table V. The models are listed as reported by the original authors to accurately reflect the experimental settings used in each work. One study does not explicitly report the underlying models employed and is therefore marked as Not Specified.
To enable consistent comparison across studies, we also categorize the evaluated language models based on their parameter count into three size categories: small, medium, and large. Small models have at most 7 billion parameters ($\leq 7B$), medium models have between 8 and 30 billion parameters, and large models exceed 30 billion parameters ($\geq 30B$). Some studies evaluate models spanning multiple size categories and are therefore labelled as mixed-scale to accurately reflect their evaluation scope. Studies that do not report model sizes or rely on proprietary systems with undisclosed back-end configurations are labelled as not specified. This classification enables a clear and reproducible analysis of model usage trends while accounting for differences in computational scale.
IV-B2 Model Openness
@@
Table IX summarizes the categorization of the selected studies based on model openness. It can be seen that the studies are evenly distributed across open-source, closed-source, and mixed model categories, with six studies in each group. In addition, although one study [33] discusses the use of proprietary models, it does not clearly specify which models are evaluated and is therefore categorized as not specified.
IV-B3 Model Domain Specialization
To characterize the domain focus of the evaluated models, we classify them as general-purpose, code-specialized, or mixed (general and code-specialized) language models. General-purpose LLMs [133] are designed for a broad range of natural language tasks without domain-specific optimization. Code-specialized LLMs [9] are trained or fine-tuned for programming-related tasks, including code generation, completion, and analysis. Studies that evaluate both general-purpose and code-specialized models are classified as mixed. Studies that do not report the domain specialization of the evaluated models are labelled as not specified.
@@ IV-B3 Model Domain Specialization
General-Purpose LLMs [96], [110], [60], [25], [105] 5
@@
The distribution of studies across these domains is presented in Table X. It can be seen that only five studies focus exclusively on general-purpose language models, while the remaining studies evaluate either code-specialized models or include at least one code-specialized model in a mixed evaluation setting.
IV-C Evaluation
IV-C1 Evaluation Metrics

### v3

# Sustainable Code Generation Using Large Language Models: A Systematic Literature Review

@@ Abstract
Large Language Models (LLMs) are increasingly used in software engineering to help developers generate, complete, translate, and fix code. However, their growing use has raised concerns about environmental sustainability. Most existing studies focus on the high energy consumption and carbon emissions of model training and inference. In contrast, much less attention has been given to the sustainability of the code generated by these models. The efficiency of generated code affects the long-term environmental impact of software systems. Inefficient code can increase CPU usage, memory consumption, execution time, and overall energy use during deployment and operation. As LLM-generated code becomes more common in real-world projects, even small inefficiencies can lead to high environmental costs over time. This paper examines existing research on the sustainability of code generated by LLMs. We conduct a systematic literature review to analyze selected primary studies and investigate the extent to which LLMs are capable of producing sustainable code. In addition, we examine how sustainability is defined and measured in this context, including the metrics and evaluation strategies used to assess energy efficiency and resource usage. We also explore whether techniques such as fine-tuning and prompt engineering influence the sustainability of generated code. Through a structured analysis of the selected studies, we categorize research efforts based on their methodological approaches, evaluation practices, and experimental settings. The findings indicate that research in this area remains relatively limited and fragmented, with no widely accepted framework for measuring or benchmarking the sustainability of LLM-generated code. These observations highlight the need for clearer definitions, standardized evaluation methods, and systematic research to support environmentally friendly AI-assisted software engineering.
@@ I Introduction
Large Language Models (LLMs) are advanced AI systems trained on large corpora of data using transformer-based architectures [126, 48]. These models are capable of understanding, reasoning over, and generating human-like language, enabling their use across a wide range of applications[145]. In recent years, LLMs have gained significant attention in the software engineering domain, where they support and automate various programming tasks. One of the most prominent applications is code generation, in which an LLM produces executable source code from natural language instructions. Closely related tasks include code completion (predicting missing or subsequent lines of code), code translation (converting source code between programming languages), bug fixing (detecting and correcting programming errors), and documentation generation (producing natural language explanations of code functionality). These capabilities are now embedded in widely used AI-assisted development tools, such as GitHub Copilot [42], CodeGeeX [139], and Amazon CodeWhisperer [7], which leverage state-of-the-art code LLMs to improve developer productivity and streamline the software development process.
LLMs [90, 47] have achieved tremendous success in Code generation [51, 100]; however, there is a growing concern about the impact of software development on the environment. Training and deploying LLMs incurs significant environmental costs, including substantial CO2 emissions and water usage [33]. According to [82], training LLaMA 3.1 with 8 billion parameters generated approximately 420 tCO2e, which is equivalent to the emissions from 83 years of electricity usage by a single U.S. household. The process also consumed 2,769 kiloliters of water, roughly equal to 24.5 years of water usage by an average American, and this is only for the training phase. Once deployed, these models continue to consume energy as users interact with them. Energy usage during inference has increased rapidly [33], and total emissions depend on how frequently the model is used. For example, if ChatGPT receives 100 million queries per day [33], and each query consumes about 0.002 kWh of energy, the total daily energy consumption would be approximately 0.2 gigawatt-hours (GWh) [124].
@@
TABLE I: Research Questions
RQ Description
RQ2 What metrics or parameters are used to evaluate sustainability?
@@
The remainder of this paper is organized as follows. Section II describes the research methodology. Section III provides insights into the extracted data, while Section IV presents the analysis and results from the extracted data. Section V presents the future research directions. Section VI outlines the validity threats and limitations. Finally, Section VII presents concluding remarks.
@@ II-A Large Language Models
Architecturally, an LLM is a stack (concatenation) of Transformer blocks [126]. To pass the text data in the transformer, each input token (word) $x_{t}$ is first mapped to a continuous embedding [6] $e_{t}\;=\;E(x_{t})\in\mathbb{R}^{d}$ where $E$ is the embedding matrix and $d$ is the hidden size. A Transformer block takes the current hidden states and applies self-attention followed by a position-wise feed-forward network (with residual connections and layer normalization omitted here for brevity). This combination of self-attention and feed-forward layers, together with residual connections and normalization, forms the core building block of the Transformer architecture used in LLMs. For more information on transformer architectures and their underlying mechanisms, readers are referred to [126].
@@ III-B Research Questions
RQ1 evaluates how effectively LLMs generate sustainable code, considering factors such as energy consumption, computational overhead, and resource usage. It examines whether sustainability arises inherently from LLMs or from deliberate design choices across different domains and programming languages.
RQ2 identifies and categorizes the metrics used to assess the sustainability of LLM-generated code, including energy consumption, carbon footprint, memory usage, and execution time. This analysis helps clarify how sustainability is defined and supports more standardized evaluation methods.
RQ3 investigates the benchmarks, datasets, and tools used to evaluate the sustainability of LLM-generated code. It assesses whether these resources are specifically designed for sustainability or adapted from general-purpose benchmarks, highlighting gaps for future development.
RQ4 examines the impact of model adaptation techniques, such as prompt engineering and fine-tuning, on the sustainability of generated code. It aims to identify practices that improve efficiency while maintaining correctness.
@@ III-C Search Strategy
Guided by the RQs, we adopt an iterative search strategy that queries digital libraries using Boolean combinations of relevant keywords. The keywords are refined after each iteration to improve coverage, and the process continues until all manually identified publications are retrieved.
@@ III-C1 Search String
Our search string was structured around two main aspects: (i) LLMs, code generation, and related techniques such as fine-tuning, and (ii) sustainability and energy efficiency. Combining these aspects ensured broad coverage of studies on sustainable code generation using LLMs. The search string was refined iteratively by evaluating its ability to retrieve relevant studies. A benchmark set of thirteen (13) publications was manually compiled from sources such as Google Search, Google Scholar, and arXiv. In the initial iteration, 10 out of 13 benchmark studies were retrieved. To capture the missing studies, additional terms (e.g., “LLM-generated code,” “energy efficiency,” and “efficient code”) were incorporated. The revised search string successfully retrieved the complete benchmark set and was therefore adopted for the remainder of the study. The final search string is presented in Table III. We searched five databases, IEEE Xplore, Compendex, Inspec, Scopus, and the ACM Digital Library, and managed the retrieved publications using an Excel spreadsheet. The search process and publication counts are shown in Fig. 1.
@@
TABLE IV: Overview of the evaluated models and their parameter-scale categories.
III-D Inclusion and Exclusion Criteria
To ensure that only relevant and high-quality studies were included in this systematic literature review, we applied a predefined set of inclusion and exclusion criteria, summarized in Table II.
We included studies that used or evaluated LLMs for code generation, addressed sustainability aspects such as energy efficiency or resource usage, and reported empirical results. Studies were excluded if they did not involve LLMs, lacked sustainability analysis, focused on unrelated tasks, or did not have accessible full text in English. The criteria were applied at multiple stages, including initial filtering, full-text screening, and snowballing, which resulted in the addition of 3 more primary studies.
Although several studies claim efficiency improvements, many do not provide empirical evaluation of energy consumption. Since our objective is to assess the sustainability of LLM-generated code based on measurable evidence, we refined the exclusion criteria to remove studies lacking energy-related analysis. As a result, only studies with explicit energy evaluation were included, reducing the final set to 18 studies. To ensure comprehensive coverage of this emerging area, we also included relevant preprints from arXiv, as sustainability-focused research on LLM-generated code is still evolving and may not yet be fully represented in peer-reviewed venues.
@@ III-E Study Selection Procedure
We then applied the refined inclusion criteria to all retrieved publications, identifying 170 studies for full-text review. To ensure consistency, the same pilot process with Fleiss’ Kappa measurement was used at this stage. After full-text screening, only 15 studies met the criteria and were included as primary studies. No additional quality assessment filters were applied, as the final pool was already small and the research area is still emerging. This approach ensured inclusiveness while maintaining relevance to our research focus.
We conducted a snowballing [64] search to identify additional primary studies. This involved reviewing both reference lists (backward snowballing) and citing papers (forward snowballing). Both approaches were applied to the 15 selected studies, using Google Scholar to retrieve citations. The same selection criteria were applied to all references and citations. This process identified 3 additional studies, bringing the final set to 18 primary studies. Fig. 1 illustrates how the initial 2,015 publications were reduced to these 18 studies used for data extraction.
TABLE V: Data Extraction Form
Category Data Items Description
Code-Related Task The primary coding task addressed (e.g., code generation, completion, etc.)
Model Adaptation Fine-Tuning Details Indicates whether the study is performing sustainability-aware fine-tuning.
IV Data Extraction Form
We created a structured data extraction form to collect information from each primary study consistently and accurately, focusing on our research questions (RQ1–RQ4). First, three researchers prepared a draft extraction form based on a small sample of studies. We then met to combine these drafts into the final extraction form shown in Table V. The categories in the form were designed to capture the main differences found in research on sustainable code generation using LLMs.
@@ IV Data Extraction Form
V Data Extraction
We present the analysis of data from the final set of 18 primary studies. This section outlines the extracted data, as summarized in Table V, providing a structured basis for analysis and enabling comparisons across all studies. This information is used in Section VI to synthesize findings and address RQ1–RQ4.
V-A General
V-A1 Context of Study
We first categorized the selected studies based on the research environment, distinguishing between academic and industrial settings. A study was labelled as academic if all authors were affiliated with universities or research institutions, and as industrial if at least one author had an industry affiliation. The distribution is summarized in Table. VI. Of the 18 studies, 13 were academic and 5 were industrial.
@@ V-A1 Context of Study
V-A2 Code-Related Task
We categorized the selected studies based on the code-related tasks they addressed. Code-related tasks refer to code generation, code completion, code optimization, and code refactoring. As several studies addressed more than one task (e.g., combining code generation with optimization or refactoring), a multi-label classification approach was adopted; therefore, a single study may appear in multiple task categories.
@@ V-A2 Code-Related Task
V-A3 Code Application Domain
@@
General-Purpose Software Development [92], [105], [122], [21], [26], [24], [125], [124], [62], [97], [10], [93], [111], [28], [11], [9] 16
@@
V-B Model
V-B1 Evaluated Models and Model Size Categories
We first present the list of models considered in each study in Table IV. The models are listed as reported by the original authors to accurately reflect the experimental settings used in each work.
To enable consistent comparison, we categorize models by parameter count into three groups: small ($\leq 7B$), medium (8–30B), and large ($\geq 30B$). Some studies evaluate models spanning multiple size categories and are therefore labelled as mixed-scale to accurately reflect their evaluation scope. Studies that do not report model sizes or rely on proprietary systems with undisclosed back-end configurations are labelled as not specified. This classification enables a clear and reproducible analysis of model usage trends while accounting for differences in computational scale.
V-B2 Model Openness
@@
Table IX summarizes the categorization of the selected studies based on model openness. It can be seen that the studies are distributed across open-source, closed-source, and mixed model categories, with five studies using open-source models and six studies each in the closed-source and mixed categories. In addition, although one study [32] discusses the use of proprietary models, it does not clearly specify which models are evaluated and is therefore categorized as not specified.
V-B3 Model Domain Specialization
To characterize the domain focus of the evaluated models, we classify them as general-purpose, code-specialized, or mixed (general and code-specialized) language models. General-purpose LLMs [128] are designed for a broad range of natural language tasks without domain-specific optimization. Code-specialized LLMs [8] are trained or fine-tuned for programming-related tasks, including code generation, completion, and analysis. Studies that evaluate both general-purpose and code-specialized models are classified as mixed. Studies that do not report the domain specialization of the evaluated models are labeled as not specified.
@@ V-B3 Model Domain Specialization
General-Purpose LLMs [92], [105], [24], [101] 4
@@
The distribution of studies across these domains is presented in Table X. It can be seen that only four studies focus exclusively on general-purpose language models, while the remaining studies evaluate either code-specialized models or include at least one code-specialized model in a mixed evaluation setting.
V-C Evaluation
V-C1 Evaluation Metrics